### v2

@@ Abstract
This study discusses the critical issues of Virtual Try-On in contemporary e-commerce and the prospective metaverse, emphasizing the challenges of preserving intricate texture details and distinctive features of the target person and the clothes in various scenarios, such as clothing texture and identity characteristics like tattoos or accessories. In addition to the fidelity of the synthesized images, the efficiency of the synthesis process presents a significant hurdle. Various existing approaches are explored, highlighting the limitations and unresolved aspects, e.g., identity information omission, uncontrollable artifacts, and low synthesis speed. It then proposes a novel diffusion-based solution that addresses garment texture preservation and user identity retention during virtual try-on. The proposed network comprises two primary modules - a warping module aligning clothing with individual features and a try-on module refining the attire and generating missing parts integrated with a mask-aware post-processing technique ensuring the integrity of the individual’s identity. It demonstrates impressive results, surpassing the state-of-the-art in speed by nearly 20 times during inference, with superior fidelity in qualitative assessments. Quantitative evaluations confirm comparable performance with the recent SOTA method on the VITON-HD and Dresscode datasets.
@@
Figure 1: Visualization for identity preservation and detail preservation compared with DCV-VTON [6]. DCI-VTON degrades clothes’ texture, struggles with symbols and characters on clothes, and alters tattoos, creating artifacts, while our approach maintains texture and tattoo fidelity.
@@ 1 Introduction
Our network comprises two primary modules - a warping module and a try-on module, integrated with post-processing blocks. The warping module is pivotal in aligning clothing with the individual’s features. It considers clothing specifics and person-related information, encompassing key points, dense pose images, and garment type-specific regions of interest (e.g., upper, lower, or full dresses). Subsequently, the try-on module refines the warped attire from the initial module, generating the missing parts of the image. The image then undergoes a conditional post-processing named mask-awareness technique to ensure the fundamental integrity of the individual’s identity. Examples of our impressive results compared to the ones from the SOTA paper [6] are depicted in detail in Fig. 1.
@@ 2.1 Virtual Try-on GAN-based Models
(GANs), have demonstrated notable potential. Most existing virtual try-on methods follow a two-stage process [12, 4, 21]. The first stage involves a warping module responsible for predicting the appearance flow for the global [12, 4], or local parts [21] of the garment to fit the target person’s pose. This is followed by a second stage, where a GAN-based generator seamlessly integrates the warped garment into the model. Although this method has shown some effectiveness, its heavy reliance on warping quality in the initial stage often leads to less-than-ideal outcomes. This is particularly evident in the realism of garment-skin boundaries and the overall try-on effect in the garment area. Despite their widespread use, these methods have not seen significant innovation to overcome these specific challenges. Therefore, we proposed an approach that aims to push the boundaries of the field by investigating the use of diffusion-based generators.
@@ 2.2 Virtual Try-on Diffusion Models
Other strategies combine the explicit warping module from GAN-based methods with a diffusion model to merge warped clothing and person images. LaDI-VTON [14] and DCI-VTON [6] leverage the capabilities of Stable Diffusion to preserve the texture and details of in-shop garments, achieving high-quality images to a certain level. However, these methods often create unintended artifacts in the final images, which remain challenging to control. Moreover, despite the outstanding generation quality of the diffusion approach, the models’ inference times are excessively long. The major reason lies in the large number of denoising steps (a large number of time steps) during the inference process.
@@ 3 Methodology
In the virtual try-on task, given an image of a person $I_{p}$ and an image of a garment $I_{g}$, we want to obtain the try-on image that portrays the person wearing the garment. The overall architecture of our approach is depicted in Figure 2. We first pre-process the person image to obtain reference information, including human parsing, 2D pose key points $J_{p}$, and dense pose $I_{dp}$. Following this, our wrapping module utilizes the reference information and produces predicted multi-scale warped-based parsing {$S_{i}$} including both the warped cloth mask $S^{i}_{m}$ and the segmentation of visible body parts $S^{i}_{bp}$, along with multi-scale appearance flows {$f_{i}$}  where the value of $i$ is in range of $[1,7]$ in resolution 512 x 384 and in range of $[1,6]$ in resolution 256 x 192. At the highest resolution level (considered here as $i=1$), these outputs are integrated with the garment and person images, in addition to preserved person information ($M_{p}$), to construct the conditional input ($I_{c}$). The conditional input $I_{c}$ in conjunction with dense pose $I_{dp}$, predicted body part segmentation $S^{i}_{bp}$, garment image $I_{g}$, and person image $I_{p}$ are fed into the try-on module to produce the final image $I_{Out}$.
@@ 3.1 Preprocessing
The procedure begins with generating human parsing maps from the person image ($I_{p}$) by employing advanced human parsing methods [13]. Then, we apply 2D key points [3], and dense pose estimation [7]; their outputs are denoted as $J_{p}$ and $I_{dp}$, respectively. Subsequently, depending on the garment type, we identify and exclude mutable sections based on the human parsing map. In detail, for dress outfits, the changeable parts encompass the entire arm (excluding the hand), the leg, and the exposed body area, including the neck, along with the area unoccupied by the clothing. For upper clothes, this changeable region now combines the arms and torso area of both the person and the clothes. Meanwhile, for lower clothes, the changeable parts are limited to the legs and lower clothing. The remaining segments are combined to create the preserved mask ($M_{p}$).
@@ 3.2 Warping Module
The warping module is the first critical component of our system, designed to handle the task of aligning an in-shop garment image ($I_{g}$) with the corresponding regions on a person image ($I_{p}$). This module is tasked with estimating the warping flow for the garment and identifying the segmentation regions for both the clothing and any occluded body parts. Our warping module adopts almost all the concepts from GP-VTON [21] except internal enhancement. Instead of estimating the local flows for certain parts of the cloth, we target the direct prediction of global flow for warping intact garments likes [12, 9, 8, 5], as depicted in Fig.3 (A). A significant enhancement in our approach is the integration of multi-head cross-attention mechanisms [18] for improved feature integration, which in turn boosts the fidelity of the warping results and overall model efficacy.
$\centering Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d}}V)\ \@add@centering$ (1)
@@
where $Q\in\mathbb{R}^{M\times d},K\in\mathbb{R}^{N\times d},V\in\mathbb{R}^{N\times d}$ represent stacked vectors of query, key, and value, respectively. Here, $M$ is the number of query vectors, $N$ is the number of key and value vectors, and $d$ is the dimension of the vector. In our setup, $Q$ represents the flattened feature of the warped garment $g^{w}_{i}$ the result of warping garment feature $g_{i}$ with input flow $f^{FF}_{in}$, while $K$ and $V$ correspond to the flattened features of the person $p_{i}$. The dot-product-based attention map $\frac{QK^{T}}{\sqrt{d}}V$ serves as an additional feature indicating the similarity between the person and the warped garment. This operation is strategically placed on the Fine Flow Block (FF-B) in the Fusion Block, as depicted in Fig. 3 (B), replacing the concatenation at the feature resolution below 64 x 48 for parameter efficiency. Our warping module detail will be depicted more in the Appendix.
@@ 3.3 Try-on Module
The objective function is then applied to the integrated output and the ground truth image. In the optimization process, our model leverages global and local conditions to generate the corresponding inferred person image. In contrast to prior studies related to the diffusion model, where the amount of noise added to the image is adjusted through a time variable, in this study, we use a fixed amount of noise, rendering the time variable redundant during the training process. At this point, our approach could be considered a single-step diffusion model. The adding noise process follows the equation:
@@
where $\alpha_{n}$ is the noise ratio and $\epsilon\sim\mathcal{N}(0,I)$. Therefore, our framework is built upon the baseline model of the Denoising Diffusion Implicit Models (DDIM) study [17], with the temporal embedding block supplanted by the CLIP feature.
@@
In the absence of the ground truth image during inference, it becomes imperative to generate an alternative conditional input to facilitate the inference pipeline. In the inference process, the local condition $I_{c}$ is constructed and sent to a post-processing block after the warping module. To be specific, the output of the first post-processing block $I^{\prime}_{c}$ is from combining the local condition image $I_{c}$ with the remaining body part extracted from the ground truth image $I_{p}$. Subsequently, it is made noisy to obtain another noise condition image $I^{\prime}_{n}$. Following a similar diffusion process as outlined in the training phase, the output image $I_{Out}$ is advanced into another post-processing block, culminating in the final result $I^{\prime}_{Out}$. This sequence ensures the generation of a refined output image, compensating for the absence of ground truth data through strategic conditioning and processing stages.
@@
The unconditional block involves obtaining body segmentation masks $S^{i}_{bp}$ here $i=1$ denoting the largest segmentation map extracted from the warping module and corresponding body masks $S_{bp}$ from the reference person image. Notation that two of these maps have the same resolution. The overlapping mask $S^{\prime}_{bp}$ is used to extract unchanged elements from $I_{p}$, which are then added to $I_{c}$ to get the coarse try-on image $I^{\prime}_{c}$. This image will be then refined by our diffusion model. The conditional post-processing block is used to re-apply the unchanged parts from the input image $I_{p}$ to the refined try-on image $I_{Out}$, to correct any undesirable modification caused by the diffusion model. It has the same structure as the unconditional one, except that it is only applied when the overlapping ratio between $S_{bp}$ and $S^{i}_{bp}$ where $i=1$, is greater than a threshold. Empirically, we found this threshold $R_{overlap}$ needs to be greater than 0.8.
$\centering R_{overlap}=\frac{\left|S^{(1)}_{bp}\bigcap S_{bp}\right|}{S^{(1)}_{bp}}=\frac{S^{\prime}_{bp}}{S^{(1)}_{bp}}\@add@centering$ (3)
Note that we apply this to every single parsing part, including the left/right hand, left/right leg, and the center body part (neck and belly), illustrated in Figure 4 (B) and (C).
@@
In terms of optimization processing for the try-on module, we utilize 2 reconstruction losses which are $\ell_{1}$ loss - $L^{tryon}_{1}$ and perceptual loss - $L^{tryon}_{per}$. Furthermore, an adversarial loss from [11] is also used for better quality results - $L^{tryon}_{adv}$. The total loss for the try-on module can be formulated as:
$\centering L^{tryon}=L^{tryon}_{1}+\alpha^{tryon}_{per}L^{tryon}_{per}+\alpha^{tryon}_{adv}L^{tryon}_{adv}\@add@centering$ (4)
where, $\alpha^{tryon}_{per}$ and $\alpha^{tryon}_{adv}$ are the balance coefficients.
@@ 4.1 Experiments Setting
Our experiments primarily use the VITON-HD dataset [4], comprising 13,679 frontal-view woman and upper clothes image pairs at 1024x768 resolution. Following prior work [4, 12], we split the dataset into a training set of 11,647 pairs and a test set of 2,032 pairs. Experiments are conducted at various resolutions, and we also assess the model on the DressCode dataset [15] for added complexity. Much like VITON-HD [4], the DressCode dataset [15] is a repository of high-quality try-on data pairs, comprising three distinct sub-datasets: dresses, upper-body, and lower-body. In total, the dataset encompasses 53,795 image pairs, distributed across 15,366 pairs for upper-body attire, 8,951 pairs for lower-body clothing, and 29,478 pairs for dresses. To maintain consistency, we apply the same human parsing and key points pose estimation methods used on the DressCode dataset [15] to the VITON-HD dataset [4].

### v3

@@ Abstract
This study discusses the critical issues of Virtual Try-On in contemporary e-commerce and the prospective metaverse, emphasizing the challenges of preserving intricate texture details and distinctive features of the target person and the clothes in various scenarios, such as clothing texture and identity characteristics like tattoos or accessories. In addition to the fidelity of the synthesized images, the efficiency of the synthesis process presents a significant hurdle. Various existing approaches are explored, highlighting the limitations and unresolved aspects, e.g., identity information omission, uncontrollable artifacts, and low synthesis speed. It then proposes a novel diffusion-based solution that addresses garment texture preservation and user identity retention during virtual try-on. The proposed network comprises two primary modules - a warping module aligning clothing with individual features and a try-on module refining the attire and generating missing parts integrated with a mask-aware post-processing technique ensuring the integrity of the individual’s identity. It demonstrates impressive results, surpassing the state-of-the-art in speed by nearly 20 times during inference, with superior fidelity in qualitative assessments. Quantitative evaluations confirm comparable performance with the recent SOTA method on the VITON-HD and Dresscode datasets. We named our model Fast and Identity Preservation Virtual TryON (FIP-VITON).
@@
Figure 1: Visualization for identity preservation and detail preservation compared with DCI-VTON and StableVITON [6, 14]. Both models tend to degrade the texture of clothes, struggle with maintaining symbols on garments, and produce in noticeable artifacts while our approach maintains the fidelity of both garment textures and tattoos.
@@ 1 Introduction
Our network comprises two primary modules - a warping module and a try-on module, integrated with post-processing blocks. The warping module is pivotal in aligning clothing with the individual’s features. It considers clothing specifics and person-related information, encompassing key points, dense pose images, and garment type-specific regions of interest (e.g., upper, lower, or full dresses). Subsequently, the try-on module refines the warped attire from the initial module, generating the missing parts of the image. The image then undergoes a conditional post-processing named mask-awareness technique to ensure the fundamental integrity of the individual’s identity. Examples of our impressive results compared to those from the SOTA papers [6, 14] are depicted in detail in Fig. 1.
@@ 2.1 Virtual Try-on GAN-based Models
(GANs), have demonstrated notable potential. Most existing virtual try-on methods follow a two-stage process [16, 4, 31]. The first stage involves a warping module responsible for predicting the appearance flow for the global [16, 4], or local parts [31] of the garment to fit the target person’s pose. This is followed by a second stage, where a GAN-based generator seamlessly integrates the warped garment into the model. Although this method has shown some effectiveness, its heavy reliance on warping quality in the initial stage often leads to less-than-ideal outcomes. This is particularly evident in the realism of garment-skin boundaries and the overall try-on effect in the garment area.
Despite their widespread use, these methods have not seen significant innovation to overcome these specific challenges. Therefore, we proposed an approach that aims to push the boundaries of the field by investigating the use of diffusion-based generators.
@@ 2.2 Virtual Try-on Diffusion Models
Other strategies combine the explicit warping module from GAN-based methods with a diffusion model to merge warped clothing and person images. LaDI-VTON [21], DCI-VTON [6], and StableVITON [14] leverage the capabilities of Stable Diffusion to preserve the texture and details of in-shop garments, achieving high-quality images to a certain level. However, these methods often create unintended artifacts in the final images, which remain challenging to control. Despite the outstanding generation quality of the diffusion approach, the models’ inference times are excessively long. The major reason lies in the large number of denoising steps during the inference process.
2.3 Diffusion Model Speed Up Techniques
Recent research has focused on accelerating the inference time of Diffusion model networks. Most of these methods are based on types of distillation techniques, such as [26, 19, 20]. Meanwhile, in specific terms of virtual try-on, no generalized approaches for multiple timestep diffusion have been distilled. Although [26] offers a direct training method that allows generating images with a single timestep, there is insufficient evidence that this method works well in a conditional high-resolution diffusion model, leaving space for future research. Furthermore, several studies applied a multimodal conditional GAN to reduce the number of timesteps to 4 or 2 [30, 23], speeding up the inference process of the diffusion model. Inspired by leveraging multimodal GAN [30, 23], we opt to experiment with a single-step diffusion-based approach employing a fixed noise level based on Denoising Diffusion Implicit Models (DDIM) [25].
@@ 3 Methodology
In the virtual try-on task, given an image of a person $I_{p}$ and an image of a garment $I_{g}$, we want to obtain the try-on image that portrays the person wearing the garment. The overall architecture of our approach is depicted in Figure 2. We first pre-process the person image to obtain reference information, including human parsing, 2D pose key points $J_{p}$, and dense pose $I_{dp}$. Following this, our wrapping module utilizes the reference information and produces predicted multi-scale warped-based parsing {$S_{i}$} including both the warped cloth mask $S^{i}_{m}$ and the segmentation of visible body parts $S^{i}_{bp}$, along with multi-scale appearance flows {$f_{i}$}. At the highest resolution ($i=1$), these outputs are integrated with the garment and person images, in addition to preserved person information ($M_{p}$), to construct the conditional input ($I_{c}$). The conditional input $I_{c}$ in conjunction with dense pose $I_{dp}$, predicted body part segmentation $S^{i}_{bp}$, garment image $I_{g}$, and person image $I_{p}$ are fed into the try-on module to produce the final image $I_{Out}$.
@@ 3.1 Preprocessing
The procedure begins with generating human parsing maps from the person image ($I_{p}$) by employing advanced human parsing methods [17]. Then, we apply 2D key points [3], and dense pose estimation [7]; their outputs are denoted as $J_{p}$ and $I_{dp}$, respectively. Subsequently, depending on the garment type, we identify and exclude mutable sections based on the human parsing map. Specifically, this mechanism is that if the garment is upper types, the upper regions of the body are omitted, which works the same as the lower types and dresses – the full body types. The remaining segments are combined to create the preserved mask ($M_{p}$).
@@ 3.2 Warping Module
In this section, we detail the structure of the Warping Module. As illustrated in Figure 3, our warping module draws inspiration from the flow estimation pipeline found in [16, 9, 8, 5, 31]. It comprises two pyramid feature extraction blocks, "Garment Feature Extraction" and "Person Feature Extraction" modules; cascade flow estimation. Our distinct contributions in this module will be elucidated below.
3.2.1 Pyramid Feature Extraction.
Our warping module leverages two Feature Pyramid Networks (FPN) [18] for extracting $N$ multi-scale person and garment features. The person feature extraction block receives inputs from the 2D human pose keypoints map $J_{p}$, dense pose $I_{dp}$, and the preserving region mask $M_{p}$. Conversely, the garment feature extraction block exclusively takes inputs from the intact in-shop garment $I_{g}$. Notably, we have no change compared to the FPN [18] used in [31] except the number of scales ($N$) varies on input resolution.
3.2.2 Cascade Flow Estimation.
Inspired by established methods [16, 9, 8, 5, 31], instead of estimating the local flows for certain parts of the cloth as in [31], we target the direct prediction of global flow for warping intact garments. Meanwhile, our warping module, depicted in Figure 3, adopts concepts from [31] with internal enhancement. Notably, our enhancement introduces cross-attention for additional feature integration, elevating the quality of the warping outcome and thereby enhancing the overall model performance.
Specifically, our module incorporates $N$ Fusion blocks designed to handle multi-scale flow maps and human parsing predictions. Each Fusion block is composed of a Coarse Flow Block (CF-B), a Fine Flow Block (FF-B), and a Global Parsing Block (GP-B), depicted in gray, blue, and green, respectively, Fig. 3(B).
In CF-B, the garment feature $g_{i}$ undergoes warping with incoming flow $f_{in}$ to produce $g^{w}_{i}$. The correlation operation from FlowNet2 [11] integrates it with the person feature $p_{i}$, and subsequent convolution layers estimate the corresponding flow $f^{\prime}_{CF}$. The refined coarse flows $f^{CF}_{out}$ result from adding $f^{\prime}_{CF}$ to $f_{in}$. FF-B, sharing CF-B’s architecture, treats $f^{CF}_{out}$ as the input flow $f^{FF}_{in}$. Diverging from CF-B, FF-B opts for multi-head cross-attention instead of correlation, using scaled dot-product attention [27]:
$\centering\textrm{Attention}(Q,K,V)=\textrm{softmax}(\frac{QK^{T}}{\sqrt{d}})V\ \@add@centering$ (1)
@@
where $Q\in\mathbb{R}^{M\times d},K\in\mathbb{R}^{N\times d},V\in\mathbb{R}^{N\times d}$ represent stacked vectors of query, key, and value, respectively. $M$ is the number of query vectors, $N$ is the number of key and value vectors, and $d$ is the dimension of the vector. In our setup, $Q$ represents the flattened feature of the warped garment $g^{w}_{i}$, while $K$ and $V$ correspond to the flattened features of the person $p_{i}$. The dot-product-based attention map $\frac{QK^{T}}{\sqrt{d}}$ serves as an additional feature indicating the similarity between the person and the warped garment. Moreover, we restrict the application of multi-head cross attention to the feature resolution below 64x48 for parameter efficiency, switching to concatenation at larger resolutions. The output of this operation is directed to a group convolution block for estimating flow $f^{\prime}_{FF}$, which is added to $f^{FF}_{in}$ to yield the fine flow $f_{out}$.
Within the Global Parsing Block (GP-B), leveraging the enhanced flow $f^{GP}_{in}$, the garment feature $g_{i}$ is warped. The newly warped feature $g^{w}_{i}$ undergoes fusion with the incoming person feature $p_{i}$ through convolution operations. The concatenated feature ${gp}_{i}$ is then processed by convolutional layers to estimate the global parsing $S_{i}$, covering background, cloth, left/right arms, center body parts (including neck and belly), and left/right legs.
3.2.3 Objective Function.
Similar to numerous prior studies employing flow-based warping models [16, 9, 8, 5, 31], our warping model follows a similar structure, incorporating a combination of various loss functions. In the training of the warping module, we utilized $\ell_{1}$ loss - $L^{w}_{1}$ and perceptual loss [12] - $L^{w}_{per}$ for the warped result. Additionally, pixel-wise cross-entropy loss $L_{ce}$, and $\ell_{1}$ loss $L^{w}_{m}$ are applied to the entire estimated parsing. Adversarial loss [13] - $L^{w}_{adv}$ is employed for both overall parsing and the warped result. Given the appearance flow’s high degree of freedom, we also include total-variation (TV) loss $L_{TV}$ as proposed in [6], effectively addressing the smoothness of the final warping result. In alignment with the approach outlined in [5], we augment a second-order smooth constraint $L_{sec}$. The total loss function for our warping module can be formulated as:
$\centering L^{w}=L^{W}_{1}+\alpha^{w}_{per}L^{w}_{per}+\alpha_{ce}L_{ce}+\alpha^{w}_{m}L^{w}_{m}+\alpha^{w}_{adv}L^{w}_{adv}+\alpha_{TV}L_{TV}+\alpha_{sec}L_{sec},\@add@centering$ (2)
with $\alpha^{w}_{per}$, $\alpha_{ce}$, $\alpha^{w}_{m}$, $\alpha^{w}_{adv}$, $\alpha_{TV}$, and $\alpha_{sec}$ are hyper-parameters controlling the contribution of each loss term to the overall loss.
@@ 3.3 Try-on Module
The objective function is then applied to the integrated output and the ground truth image. Our model leverages global and local conditions in the optimization process to generate the corresponding inferred person image. In contrast to prior studies related to the diffusion model, where the amount of noise added to the image is adjusted through a time variable, in this study, we use a fixed amount of noise, rendering the time variable redundant during the training process. At this point, our approach could be considered a single-step diffusion model. The adding noise process follows the equation:
@@
where $\alpha_{n}$ is the noise ratio and $\epsilon\sim\mathcal{N}(0,I)$. Therefore, our framework is built upon the baseline model of the DDIM study [25], with the temporal embedding block supplanted by the CLIP feature.
@@
In the absence of the ground truth image during inference, it becomes imperative to generate an alternative conditional input to facilitate the inference pipeline. In the inference process, the local condition $I_{c}$ is constructed and sent to a post-processing block after the warping module. Specifically, the output of the first post-processing block $I^{\prime}_{c}$ is created by combining the local condition image $I_{c}$ with the remaining body parts extracted from the ground truth image $I_{p}$. This combined image is then made noisy to obtain another noise-conditioned image $I^{\prime}_{n}$. Following a similar diffusion process as outlined in the training phase, the output image $I_{Out}$ is advanced into another post-processing block, culminating in the final result $I^{\prime}_{Out}$. This sequence ensures the generation of a refined output image, compensating for the absence of ground truth data through strategic conditioning and processing stages.
@@
The unconditional post-processing block involves obtaining body segmentation masks $S^{i}_{bp}$ here $i=1$ denoting the largest segmentation map extracted from the warping module and corresponding body masks $S_{bp}$ from the reference person image. The overlapping mask $S^{\prime}_{bp}$ is used to extract unchanged elements from $I_{p}$, which are then added to $I_{c}$ to get the coarse try-on image $I^{\prime}_{c}$.
The conditional post-processing block is used to re-apply the unchanged parts from the input image $I_{p}$ to the refined try-on image $I_{Out}$, correcting any undesirable modifications caused by the diffusion model. It has the same structure as the unconditional one, except that it is only applied when the overlapping ratio between $S_{bp}$ and $S^{i}_{bp}$ where $i=1$, is greater than a threshold. Empirically, we found this threshold $R_{\textrm{overlap}}$ needs to be greater than 0.8. This experiment is detailed in the Appendix, providing the applied rate of each specific threshold.
$\centering R_{\textrm{overlap}}=\frac{\left|S^{(1)}_{bp}\bigcap S_{bp}\right|}{S^{(1)}_{bp}}=\frac{S^{\prime}_{bp}}{S^{(1)}_{bp}}\@add@centering$ (4)
Notably, we apply this to every single parsing part, including the left/right hand, left/right leg, and center body part (neck and belly), shown in Fig. 4(B),(C).
@@
In terms of optimization processing for the try-on module, we utilize 2 reconstruction losses which are $\ell_{1}$ loss - $L^{\textrm{tryon}}_{1}$ and VGG perceptual loss [12] - $L^{\textrm{tryon}}_{per}$. Furthermore, an adversarial loss from [13] is also used for better quality results - $L^{\textrm{tryon}}_{adv}$. The total loss for the try-on module can be formulated:
$\centering L^{\textrm{tryon}}=L^{\textrm{tryon}}_{1}+\alpha^{\textrm{tryon}}_{per}L^{\textrm{tryon}}_{per}+\alpha^{\textrm{tryon}}_{adv}L^{\textrm{tryon}}_{adv}\@add@centering$ (5)
where, $\alpha^{\textrm{tryon}}_{per}$ and $\alpha^{\textrm{tryon}}_{adv}$ are the balance coefficients.
@@ 4.1 Experiments Setting
Our experiments primarily use the VITON-HD dataset [4], comprising 13,679 frontal-view woman and upper clothes image pairs at 1024x768 resolution. Following prior work [4, 16], we split the dataset into a training set of 11,647 pairs and a test set of 2,032 pairs. Experiments are conducted at various resolutions, and we also assess the model on the DressCode dataset [22] for added complexity. Much like VITON-HD [4], the DressCode dataset [22] is a repository of high-quality try-on data pairs, comprising three distinct sub-datasets: dresses, upper-body, and lower-body. In total, the dataset encompasses 53,795 image pairs, distributed across 15,366 pairs for upper-body attire, 8,951 pairs for lower-body clothing, and 29,478 pairs for dresses. To maintain consistency, we apply the same human parsing and key points pose estimation methods used on the DressCode [22] to the VITON-HD [4].